A stock technical analyst uses smoothing methods to predict stock prices. The following table contains the monthly price information for Walmart.

Date Price
11/30/2018 95.62
12/31/2018 91.71
1/31/2019 94.35
2/28/2019 97.46
3/29/2019 96.54
4/30/2019 101.80
5/31/2019 100.94
6/28/2019 109.95
7/31/2019 109.84
8/30/2019 114.26
9/30/2019 118.68
10/31/2019 117.26

What is the predicted share price for August 30, 2019 if this technical analyst uses an exponential smoothing method with a smoothing constant of .7 and a 5-month moving average as the initial forecast share price for April 30, 2019?
Answer #1

Exponential smoothing method is given by the following formula:

S0 = X0,

S4 = Q * X+ + (1 - a) * St-1

In our case, initial forecast share price for April 30, 2019 is assumed as 5 month moving average i.e. average of last 5 months prices.

Hence forecasted share price for April will be SApril = 0.7*101.8 + (1-0.7) * (95.62 + 91.71 + + 94.35 + 97.46 + 96.54)/5 = 99.8008

Now use this forecasted price to calculate the next months price as follows"

SMay = 0.7*XMay+ (1-0.7) * SApril  

SMay = 0.7*100.94+ (1-0.7) * 99.8008 = 100.59824

SJune = 0.7*XJune+ (1-0.7) * SMay  

SJune = 0.7*109.95+ (1-0.7) * 100.59824= 107.144472

SJuly = 0.7*XJuly+ (1-0.7) * SJune  

SJuly = 0.7*109.84+ (1-0.7) * 107.144472= 109.0313416

SAug = 0.7*XAug+ (1-0.7) * SJuly  

SAug = 0.7*114.26+ (1-0.7) * 109.0313416= 112.6914025

Answer: Predicted share price for August 30, 2019 is $112.6914025